# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ I Introduction
In the last decade, experimental results brought compelling evidence that in ultra-relativistic heavy-ion collisions, a plasma of quarks and gluons (QGP) is produced STAR:2011fbd ; PHENIX:2006dpn ; CMS:2016xef ; Matsui:1986dk . Such a state of matter has earlier been predicted by lattice gauge calculations (lQCD). Modern lQCD calculations robustly predict a crossover transition between the hadronic and the QGP phases, which happens (for vanishing chemical potential) approximately in the temperature range $T=154\pm 9$ MeV Borsanyi:2013bia ; Bazavov:2014pvz . This new QGP state has, however, a very short lifetime of the order of a couple of fm$/c$ before it transforms into hadrons which are finally observed in the detectors. The experimental and theoretical study of the properties of the QGP is presently one of the major challenges in hadron physics. Since quarks and gluons cannot be observed directly, it is necessary to study them by the analysis of hadron observables and it turns out that the multiplicity of the majority of hadrons, those which contain only up, down and strange quarks is well described by the assumption that the expanding system is at hadronization in statistical equilibrium Andronic:2005yp . More information can be obtained from the detailed analysis of the results of transport approaches, which describe the time evolution of the QGP. Employing the technique of Bayesian analysis constraints on the shape of the initial state, shear Bernhard:2016tnd and bulk Bernhard:2019bmu viscosity of the QGP medium and its equation of state Pratt:2015zsa can be established.
@@
The first steps towards such a study have been paved by Wiedemann Wiedemann:2000za , who showed that the BDMPS-Z approach, which is defined on a Keldysh time contour and hence difficult to translate into an transport approach because it requires the knowledge of the future development of the system, could be transformed into an opacity expansion, which can serve subsequently Zapp:2011ya as a basis for a numerical study of the LPM effect and its consequences for the energy loss of a fast particle in a static medium.
@@
It is one of the goals of ultra-relativistic heavy-ion collisions to study how fast partons, created in the interaction of the incoming protons, pass a plasma of quark and gluons, which is created in such collisions. To achieve this goal, one has to embed the results obtained for static QGP matter in a transport approach, which describes the time evolution of the heavy ion collisions. It is the purpose of this paper to outline the model, which we have developed for this. It is designed to describe the evolution of a parton in an expanding medium and we will present here all elements of this approach. This approach is a hybrid approach because the energy loss of the parton has two origins. Being produced in heavy ion collisions, the parton has initially a high virtuality, which decreases due to the emission of gluons. Such emissions take also place in the absence of the QGP but its presence modifies the energy loss of the leading parton. When the partons have reached low virtuality, the collisions with the QGP create an additional induced energy loss due to the scattering of the partons with the constituents of the QGP. This two step scenario is similar to the decomposition adopted in JETSCAPE JETSCAPE:2017eso .
In section II we present an outline of our approach, based, for its high virtuality component, on the ideas of the YaJEM Renk:2008pp approach, while its low virtuality component follows the work of Wiedemann and ZappZapp:2011ya , revisited in order to take better care of the low-energy Gunion Bertsch (GB) gluons and properly calibrated on recent theoretical estimates of the radiation yield. This is followed by a introduction to the details. We start out with a discussion of how we describe the time evolution of highly virtual partons and its modification in a QGP in section III. Then, in section IV, we present all the ingredients of the low virtuality component, including the model adopted for elastic scattering, the elementary radiation amplitude of low-virtuality partons, which scatter with the QGP partons, and how these cross sections can be modeled by a Monte Carlo approach, necessary for its use in a transport theory. The quantum interference of the radiation amplitudes, the LPM effect, and its implementation in a transport theory are discussed in subsection IV.5.
@@
In section VI we conclude our results. The application of our model to heavy ion collisions, in which the QGP is expanding, as well as the comparison of our results with experimental data, we postpone to a future publication. Throughout all this manuscript, we consider natural units with $\hbar=c=1$.
@@ II Outline of our hybrid approach
Hard scatterings between projectile and target nucleons can lead to a large momentum transfer either of the scattered partons or of the produced partons. Such partons are called jets or jet seeds and have a high virtuality, which is, due to the lack of QCD based results, differently parameterized in different approaches. Subsequently, this parton loses its virtuality by emission of other partons. The splitting into several partons is described by the Sudakov form factor. In a medium, while losing its virtuality, the parton can also scatter with the QGP constituents. An effective way to describe the early jet evolution in the QGP medium is to assume that the scattering with the medium induce a gain of virtuality, as in YaJEM Renk:2008pp . Splitting ends when the virtuality of the parton arrives at the lowest virtuality scale $2Q_{0}$ with $Q_{0}\simeq\Lambda_{\text{QCD}}$. Below $2Q_{0}$ the parton enters the low-virtuality kinetic regime. More sophisticated transition criteria will be explored in a future work.
@@
As soon as they enter the kinetic regime, the jet partons are considered as an on-shell particles with a thermal mass and scatter elastically with the partons (gluons and quarks) of the QGP, which is assumed to be in thermal equilibrium. These processes are modeled through a differential rate of transverse momentum exchange with the QGP particles evaluated by using standard QCD matrix elements where the exchanged gluon has an infrared regulator of $\mu^{2}=\kappa m_{D}^{2}$ for the Debye mass $m_{D}\sim gT$ and sampled in a Monte Carlo procedure. This allows to assure momentum and energy conservation. For the mass of the QGP particles two limiting cases are studied, $m_{\text{q}}=0$ and $m_{\text{q}}\to\infty$, the later choice being essentially considered for the purpose of benchmarking our results on theoretical calculations as BDMPS-Z.
@@
As said, in vacuum a virtual gluon with an energy $\omega$ and a transverse momentum $k_{\perp}$, created in inelastic collisions, needs the time $t=\omega/k_{\perp}^{2}$ to be formed. In a strongly interacting medium the gluon may scatter with the medium while being formed. These collisions modify the transverse momentum of the gluon and cause a change of the formation time. To be able to describe gluon emission in a medium in a wide range of gluon energies and hence of formation times we start for the inelastic collisions from the Gunion-Bertsch approach Gunion:1981qs ; Aichelin:2013mra , which describes the emission of a gluon with zero formation time. The finite formation time is then included by the rejection of the formation of a real gluon depending on the future path of the gluon in the medium. By this we can describe Bethe Heitler regime as well as both regimes of BDMPS-Z. If the parton-gluon antenna resulting from this seed does not suffer an elastic scattering before its formation time, the gluon is assumed to be formed and both partons (the projectile parton and the formed gluon) are free to interact with the QGP independently. If, on the contrary, the projectile parton scatters inelastically with another QGP parton before the gluon is formed, the emission amplitudes interfere, leading to the QCD equivalent of the Landau-Pomeranchuk-Migdal (LPM) effect. In practice we apply the heuristic BDMPS procedure to count the $N_{s}$ scattering centers, on which the virtual gluon scatters before being formed, and accept the formed gluon then with the probability 1/$N_{s}$, means effectively we replace the $N_{s}$ scattering centers a single one. If, finally, the formation time is even larger than the size of the system, the virtual gluon cannot be formed. As mentioned in the introduction, these three domains exist already when a fast parton traverses a brick of QCD matter and it is difficult to describe them with causal Monte Carlo approaches because the formation time is not known at the moment when the virtual gluon is formed. The approach, which we present here, is however more ambitious. We want to describe finally the gluon creation in a rapidly expanding QGP, as created in ultra-relativistic heavy-ion collisions, using the seed-rejection algorithm presented in details below.
@@
In this article, we focus on the basic benchmark of the energy loss, which a high energy parton suffers in QGP matter. Therefore, even though the framework is constructed for arbitrary space-time temperature profile, for the purpose of benchmarking and comparison to known semi-analytical results, we approximate the QGP medium by homogeneous bricks of temperature $T=400$ MeV and fixed lengths, $L=1$, $2$, and $8$ fm. We perform the calculation with a mono-energetic stream of partons with an initial energy of $E=100$ GeV.
@@ III Numerical realization of our approach: the high-virtuality component
In the high virtuality regime, jets evolve as virtuality-ordered parton cascades from an initial maximal virtuality scale $Q_{\uparrow}$ down to a minimal virtuality scale $Q_{\rm min}$. While the maximal initial virtuality of the parton is $Q_{\uparrow}=E$, the real initial virtuality $Q_{1}$ is obtained by sampling the Sudakov factor, between $Q_{\rm min}$ and $Q_{\uparrow}$.  [^3] Therefore our initial virtuality is different from the calculations of the JETSCAPE collaboration JETSCAPE:2019udz which start out with $Q^{2}=p_{T}^{2}/2$. Quark jets, e.g., are assumed to evolve by emission of gluons (bremsstrahlung) as in the vacuum, however with modifications induced by the medium through which they travel. In the next subsection we first focus on the DGLAP algorithm adopted to describe such a jet evolution in the vacuum. We take $Q_{\rm min}=2Q_{0}$, with $Q_{0}=0.3$ GeV $\simeq\Lambda_{\text{QCD}}$. In the subsequent subsection we introduce the medium modifications.
@@ III.1 Evolution in vacuum
    Both the virtualities $Q_{b}$ and $Q_{c}$ of the produced partons $b$ and $c$ are drawn randomly from the partition functions $S_{b}(Q_{a},Q_{b})$ and $S_{c}(Q_{a},Q_{c})$, respectively. In other terms, both $Q_{\uparrow c}$ and $Q_{\uparrow b}$ in eq. (3) are taken as $Q_{a}$. This is done within a rejection loop: For a given value of $Q_{a}$, first $Q_{b}$ and then $Q_{c}$ are selected. If $Q_{a}^{2}\geq Q_{b}^{2}+Q_{c}^{2}$, the values are accepted, otherwise both $Q_{b}$ and $Q_{c}$ are drawn again. The reason for this rejection loop is to guarantee values of $Q_{b}$ and $Q_{c}$ that allow for parton momenta that are kinematically possible.555 This rejection loop is not strictly necessary as the rejection can also be done at the very end of the selection of the splitting. However selecting already here virtualities that allow for kinematically possible parton momenta makes the whole algorithm more effective.
@@
Figure 1: Distribution of the ratio of the virtuality $Q_{1}$, the staring point of the jet evolution and the initial maximum virtuality $Q_{\uparrow}=E_{\text{ini}}$, averaged over all simulated parton cascades, for parton cascades with $E_{\text{ini}}=10$ and $50$ GeV.
@@
It is interesting to see the time evolution of the jets in this high virtuality regime. Fig. 2 shows the time evolution of the squared virtuality $\langle Q^{2}_{\text{tagged}}\rangle$ of the initially tagged quark for parton cascades with an initial quark and with initial energies of $E_{\text{ini}}=10$ and $50$ GeV. For this we follow the virtuality of the fermion at each splitting. The virtuality is averaged over $10^{6}$ simulated cascades. The average parton virtuality decreases fast, in agreement with the assumption of a time-like parton cascade with a strong ordering in the parton virtualities, The higher the initial energy is, the faster the initial virtuality decreases and, due to the increasing number of splittings, the longer it takes until the quark becomes on-shell. The mean virtuality of both jets decreases in the same pace for intermediate times. At $t=1$ fm$/c$ the virtuality is reduced by more than one order of magnitude, independent of the initial energy. At $t=5$ fm$/c$, jets with an initial virtuality of $50$ GeV finally arrive at a virtuality of the order of $Q_{0}$. [^7] This shows that the jet evolution in our modeling is not governed by a single time scale, even for its evolution in vacuum. In the medium the decrease of the virtuality is slower at later times because the leading partons gain virtuality in collisions with the medium.
@@
Fig. 3 shows the average of the sum of virtualites squared of all partons in the cascade of a jet as a function of the time. The sum of all virtualities decreases much slower as a function of time than the virtuality of the leading parton because the part of virtuality of the initial parton is transferred to daughter partons with a lower virtuality, which live longer. Hence it takes considerably longer until these daughter partons are on the mass shell as compared to the time for the leading parton. The curves are observed to flatten out at large times, when all the partons in the cascade arrive at their final virtuality of the order of $Q_{0}$. The larger total virtuality at larger times for the more energetic jet shows that more splittings have occurred and therefore more partons have been created. This is confirmed in Fig. 4 where the distribution of splittings (normalized to one) is shown.

### v2

@@ I Introduction
In the last decade, experimental results brought compelling evidence that in ultra-relativistic heavy-ion collisions, a plasma of quarks and gluons (QGP) is produced STAR:2011fbd ; PHENIX:2006dpn ; CMS:2016xef ; Matsui:1986dk . Such a state of matter has earlier been predicted by lattice gauge calculations (lQCD). Modern lQCD calculations robustly predict a crossover transition between the hadronic and the QGP phases, which happens (for vanishing chemical potential) approximately in the temperature range $T=154\pm 9$ MeV Borsanyi:2013bia ; Bazavov:2014pvz . This new QGP state has, however, a very short lifetime of the order of a couple of fm$/c$ before it transforms into hadrons, which are finally observed in the detectors. The experimental and theoretical study of the properties of the QGP is presently one of the major challenges in hadron physics. Since quarks and gluons cannot be observed directly, it is necessary to study them by the analysis of hadron observables and it turns out that the multiplicity of the majority of hadrons, those which contain only up, down and strange quarks is well described by the assumption that the expanding system is at hadronization in statistical equilibrium Andronic:2005yp . More information can be obtained from the detailed analysis of the results of transport approaches, which describe the time evolution of the QGP. Employing the technique of Bayesian analysis constraints on the shape of the initial state, shear Bernhard:2016tnd and bulk Bernhard:2019bmu viscosity of the QGP medium and its equation of state Pratt:2015zsa can be established.
@@
The first steps towards such a study have been paved by Wiedemann Wiedemann:2000za , who showed that the BDMPS-Z approach, which is defined on a Keldysh time contour and hence difficult to translate into a transport approach because it requires the knowledge of the future development of the system, could be transformed into an opacity expansion, which can serve subsequently Zapp:2011ya as a basis for a numerical study of the LPM effect and its consequences for the energy loss of a fast particle in a static medium.
@@
It is one of the goals of ultra-relativistic heavy-ion collisions to study how fast partons, created in the interaction of the incoming protons, pass a plasma of quark and gluons, which is created in such collisions. To achieve this goal, one has to embed the results obtained for static QGP matter in a transport approach, which describes the time evolution of the heavy ion collisions. It is the purpose of this paper to outline the model, which we have developed for this. It is designed to describe the evolution of a parton in an expanding medium and we will present here all elements of this approach. This approach is a hybrid approach because the energy loss of the parton has two origins. Being produced in heavy ion collisions, the parton has initially a high virtuality, which decreases due to the emission of gluons. Such vacuum-like emissions take also place in the absence of the QGP but its presence modifies the energy loss of the leading parton. The branching is considered to be virtuality ordered, similar to JEWEL Zapp:2008zz ; Zapp:2012ak and MATTERMajumder:2010qh . This allows to embed the algorithm in future easily in the space-time evolution of the medium through which the jet travels. Spin effects are not considered in the present version of the approach. When the partons have reached low virtuality, the evolution switches towards a time-ordered scheme of the BDMPS-Z type, where the collisions with the QGP create an additional induced energy loss due to the scattering of the partons with the constituents of the QGP. This two step scenario is similar to the decomposition adopted in JETSCAPE JETSCAPE:2017eso .
In section II we present an outline of our approach, based, for its high virtuality component, on the ideas of the YaJEM Renk:2008pp approach, while its low virtuality component follows the same strategy as developed in Zapp:2011ya . We consider, however, the influence of the phase space limitation in the gluon emission, what allows us to explore the whole $(\omega,k_{T})$ space. This permits as well to calculate the recoil of the QGP parton, which may serve as a source term in hydrodynamical calculations of the expanding QGP. We will show that for gluon energies larger than the critical energy $\omega_{c}$ we reproduce the GLV results Gyulassy:2000er , what was not addressed in Zapp:2011ya , which considered the harmonic approximation. Our approach allows to go beyond the eikonal limit for the kinematical variables as well as for the phase accumulation and can therefore handle the kinematical region explored in heavy ion collisions. We study furthermore the influence of finite masses of the QGP partons. Coming to the details of our approach, we start out with a discussion of how we describe the time evolution of highly virtual partons and its modification in a QGP in section III. Then, in section IV, we present all the ingredients of the low virtuality component, including the model adopted for elastic scattering, the elementary radiation amplitude of low-virtuality partons, which scatter with the QGP partons, and how these cross sections can be modeled by a Monte Carlo approach, necessary for its use in a transport theory. The quantum interference of the radiation amplitudes, the LPM effect, and its implementation in a transport theory are discussed in subsection IV.5.
@@
In section VI we conclude our results. The application of our model to heavy ion collisions, in which the QGP is expanding, as well as the comparison of our results with experimental data, we postpone to a future publication. Throughout this manuscript, we consider natural units with $\hbar=c=1$.
@@ II Outline of our hybrid approach
Hard scatterings between projectile and target nucleons can lead to a large momentum transfer either of the scattered partons or of the produced partons. Such partons are called jets or jet seeds and have a high virtuality, which is, due to the lack of QCD based results, differently parameterized in different approaches. Subsequently, this parton loses its virtuality by emission of other partons. The splitting into several partons is described by the Sudakov form factor. In a medium, while losing its virtuality, the parton can also scatter with the QGP constituents. An effective way to describe the early jet evolution in the QGP medium is to assume that the scattering with the medium induces a gain of virtuality, as in YaJEM Renk:2008pp . Whenever the transverse-momentum broadening, acquired by a parton from the QGP during its formation time $t_{f}$, is larger then its own transverse momentum at "creation", one switches the evolution from the high virtuality cascade to the low virtuality evolution. This criterion is applied to each parton individually and adopted for instance in Caucal:2018dla when discussing the transition from vacuum-like emissions to induced BDMPS-Z gluons. Since the gluon formation time writes $t_{f}\simeq\omega/k_{T}^{2}$, the criteria to switch to the low virtuality evolution is $\hat{q}\times\frac{\omega}{k_{T}^{2}}\gtrsim k_{T}^{2}\Leftrightarrow\omega^{3}\theta^{4}\lesssim\hat{q}$, where $\theta=k_{T}/\omega$ is the gluon emission angle in the splitting. In Caucal:2018dla the relation $\omega^{3}\theta^{4}\simeq\hat{q}$ indeed corresponds to the locus in the $(\omega,k_{T})$ plane describing the end of the vacuum-like cascade. As $k_{T}^{2}\approx zQ^{2}$ – see eq. (13) – , where $Q$ is the virtuality of the parent parton while $z=\omega/E$ is the energy fraction of the daughter parton, this condition is equivalent to
$\frac{zQ^{4}}{E}\lesssim\hat{q}\,,$ (3)
which we adopt in our hybrid approach. This criterion differs slightly from the one advocated in JETSCAPE JETSCAPE:2021ehl where a switching virtuality $Q_{\rm sw}^{2}=\sqrt{E\hat{q}}$ is applied. [^2] Although in practice, an average value is retained. It is also equivalent to require that the transition takes place when the formation time of the gluon, emitted at high virtuality, exceeds the formation time of the gluon emitted in the low virtuality BMDPS-Z regime. This strategy of branching towards the process which has the smallest formation time is found in other MC approaches, like JEWEL Zapp:2012ak . However, contrarily to JEWEL where the LPM effects are implemented in a modified virtuality-ordered DGLAP evolution equation incorporating formation time effects, our approach implements these LPM effects in a genuine time-ordered evolution, similarly to Zapp:2011ya . In the absence of any medium, we adopt the standard hypothesis that splitting ends when the virtuality of the parton arrives at the lowest virtuality scale $2Q_{0}$ with $Q_{0}\simeq\Lambda_{\text{QCD}}$. Below $2Q_{0}$ the parton enters the low-virtuality kinetic regime.
@@
As soon as they enter the kinetic regime, the jet partons are considered as on-shell particles with a thermal mass and scatter elastically with the partons (gluons and quarks) of the QGP, which is assumed to be in thermal equilibrium. These processes are modeled through a differential rate of transverse momentum exchange with the QGP particles evaluated by using standard QCD matrix elements where the exchanged gluon has an infrared regulator of $\mu^{2}=\kappa m_{D}^{2}$ for the Debye mass $m_{D}\sim gT$ and sampled in a Monte Carlo procedure. This allows to assure momentum and energy conservation. For the mass of the QGP particles two limiting cases are studied, $m_{\text{q}}=0$ and $m_{\text{q}}\to\infty$, the later choice being essentially considered for the purpose of benchmarking our results on theoretical calculations as BDMPS-Z.
@@
As said, in vacuum a virtual gluon with an energy $\omega$ and a transverse momentum $k_{\perp}$, created in inelastic collisions, needs the time $t=\omega/k_{\perp}^{2}$ to be formed. In a strongly interacting medium the gluon may scatter with the medium while being formed. These collisions modify the transverse momentum of the gluon and cause a change of the formation time. To be able to describe gluon emission in a medium in a wide range of gluon energies and hence of formation times we start for the inelastic collisions from the Gunion-Bertsch approach Gunion:1981qs ; Aichelin:2013mra , which describes the emission of a gluon with zero formation time. The finite formation time is then included by the rejection of the formation of a real gluon depending on the future path of the gluon in the medium. By this we can describe the Bethe Heitler regime as well as both regimes of BDMPS-Z. If the parton-gluon antenna resulting from this seed does not suffer an elastic scattering before its formation time, the gluon is assumed to be formed and both partons (the projectile parton and the formed gluon) are free to interact with the QGP independently. If, on the contrary, the projectile parton scatters inelastically with another QGP parton before the gluon is formed, the emission amplitudes interfere, leading to the QCD equivalent of the Landau-Pomeranchuk-Migdal (LPM) effect. In practice we apply the heuristic BDMPS procedure to count the $N_{s}$ scattering centers, on which the virtual gluon scatters before being formed, and accept the formed gluon then with the probability 1/$N_{s}$, means effectively we replace the $N_{s}$ scattering centers a single one. If, finally, the formation time is even larger than the size of the system, the virtual gluon cannot be formed. As mentioned in the introduction, these three domains exist already when a fast parton traverses a brick of QCD matter and it is difficult to describe them with causal Monte Carlo approaches because the formation time is not known at the moment when the virtual gluon is formed. The approach, which we present here, is however more ambitious. We want to describe finally the gluon creation in a rapidly expanding QGP, as created in ultra-relativistic heavy-ion collisions, using the seed-rejection algorithm presented in details below.
@@
In this article, we focus on the basic benchmark of the energy loss, which a high energy parton suffers in QGP matter. Therefore, even though the framework is constructed for an arbitrary space-time temperature profile, for the purpose of benchmarking and comparison to known semi-analytical results, we approximate the QGP medium by homogeneous bricks of temperature $T=400$ MeV and fixed lengths, $L=1$, $2$, and $8$ fm. We perform the calculation with a mono-energetic stream of partons with an initial energy of $E=100$ GeV.
@@ III Numerical realization of our approach: the high-virtuality component
In the high virtuality regime, jets evolve as virtuality-ordered parton cascades from an initial maximal virtuality scale $Q_{\uparrow}$ down to a minimal virtuality scale $Q_{\rm min}$. A virtuality ordering has been chosen here because 1/Q is the lifetime of the virtual state and therefore the algorithm can be easily embedded in the space time evolution of the environment. For future applications we consider to proceed to an angular ordering by limiting the kinematical range of the daughter partons Zapp:2008zz ; Rohrmoser:2017vsa .
While the maximal initial virtuality of the parton is $Q_{\uparrow}=E$, the real initial virtuality $Q_{1}$ is obtained by sampling the Sudakov factor, between $Q_{\rm min}$ and $Q_{\uparrow}$.  [^4] Therefore our initial virtuality is different from the calculations of the JETSCAPE collaboration JETSCAPE:2019udz which start out with $Q^{2}=p_{T}^{2}/2$. Quark jets, e.g., are assumed to evolve by emission of gluons (bremsstrahlung) as in the vacuum, however with modifications induced by the medium through which they travel. In the next subsection we first focus on the DGLAP algorithm adopted to describe such a jet evolution in the vacuum. We take $Q_{\rm min}=2Q_{0}$, with $Q_{0}=0.3$ GeV $\simeq\Lambda_{\text{QCD}}$. In the subsequent subsection we introduce the medium modifications. In order to maximize the effect of the medium on the high-virtuality component, we mostly keep the same value of $Q_{\rm min}$ as in the vacuum, disregarding for the time being the criteria given at eq. (3). For the ratio of the in-medium primary Lund plane and its equivalent in the vacuum, we provide, however, both results obtained by either applying this prescription – Fig. 9 – or by using the switch criteria expressed in eq. eq. (3) – Fig.9 –, in order to illustrate the consequence of such criteria on the differential gluon-spectrum.
@@ III.1 Evolution in vacuum
    Both the virtualities $Q_{b}$ and $Q_{c}$ of the produced partons $b$ and $c$ are drawn randomly from the partition functions $S_{b}(Q_{a},Q_{b})$ and $S_{c}(Q_{a},Q_{c})$, respectively. In other terms, both $Q_{\uparrow c}$ and $Q_{\uparrow b}$ in eq. (4) are taken as $Q_{a}$. This is done within a rejection loop: For a given value of $Q_{a}$, first $Q_{b}$ and then $Q_{c}$ are selected. If $Q_{a}^{2}\geq Q_{b}^{2}+Q_{c}^{2}$, the values are accepted, otherwise both $Q_{b}$ and $Q_{c}$ are drawn again. The reason for this rejection loop is to guarantee values of $Q_{b}$ and $Q_{c}$ that allow for parton momenta that are kinematically possible.666 This rejection loop is not strictly necessary as the rejection can also be done at the very end of the selection of the splitting. However selecting already here virtualities that allow for kinematically possible parton momenta makes the whole algorithm more effective.
@@
Figure 1: Distribution of the ratio of the virtuality $Q_{1}$, the starting point of the jet evolution and the initial maximum virtuality $Q_{\uparrow}=E_{\text{ini}}$, averaged over all simulated parton cascades, for parton cascades with $E_{\text{ini}}=10$ and $50$ GeV.
@@
It is interesting to see the time evolution of the jets in this high virtuality regime. Fig. 2 shows the time evolution of the squared virtuality $\langle Q^{2}_{\text{tagged}}\rangle$ of the initially tagged quark for parton cascades with an initial quark and with initial energies of $E_{\text{ini}}=10$ and $50$ GeV. For this we follow the virtuality of the fermion at each splitting. The virtuality is averaged over $10^{6}$ simulated cascades. The average parton virtuality decreases fast, in agreement with the assumption of a time-like parton cascade with a strong ordering in the parton virtualities. The higher the initial energy, the faster the initial virtuality decreases and, due to the increasing number of splittings, the longer it takes until the quark becomes on-shell. The mean virtuality for both energies decreases in the same pace for intermediate times. At $t=1$ fm$/c$ the virtuality is reduced by more than one order of magnitude, independent of the initial energy. At $t=5$ fm$/c$, particles with an initial virtuality of $50$ GeV arrive finally at a virtuality of the order of $Q_{0}$. [^8] This shows that the jet evolution in our modeling is not governed by a single time scale, even for its evolution in vacuum. In the medium the decrease of the virtuality is slower at later times because the leading partons gain virtuality in collisions with the medium.
@@
Fig. 3 shows the average of the sum of virtualities squared of all partons in the cascade of a jet as a function of the time. The sum of all virtualities decreases much slower as a function of time than the virtuality of the leading parton because part of the virtuality of the initial parton is transferred to daughter partons with a lower virtuality, which live longer. Hence it takes considerably longer until these daughter partons are on the mass shell as compared to the time for the leading parton. The curves are observed to flatten out at large times, when all the partons in the cascade arrive at their final virtuality of the order of $Q_{0}$. The larger total virtuality at larger times for the more energetic jet shows that more splittings have occurred and therefore more partons have been created. This is confirmed in Fig. 4 where the distribution of splittings (normalized to one) is shown.